# Text-to-CAD Generation Through Infusing Visual Feedback in Large Language Models

### v2

@@ Abstract
Creating Computer-Aided Design (CAD) models requires significant expertise and effort. Text-to-CAD, which converts textual descriptions into CAD parametric sequences, is crucial in streamlining this process. Recent studies have utilized ground-truth parametric sequences, known as sequential signals, as supervision to achieve this goal. However, CAD models are inherently multimodal, comprising parametric sequences and corresponding rendered visual objects. Besides, the rendering process from parametric sequences to visual objects is many-to-one. Therefore, both sequential and visual signals are critical for effective training. In this work, we introduce CADFusion, a framework that uses Large Language Models (LLMs) as the backbone and alternates between two training stages: the sequential learning (SL) stage and the visual feedback (VF) stage. In the SL stage, we train LLMs using ground-truth parametric sequences, enabling the generation of logically coherent parametric sequences. In the VF stage, we reward parametric sequences that render into visually preferred objects and penalize those that do not, allowing LLMs to learn how rendered visual objects are perceived and evaluated. These two stages alternate throughout the training, ensuring balanced learning and preserving benefits of both signals. Experiments demonstrate that CADFusion significantly improves performance, both qualitatively and quantitatively.
@@ 1 Introduction
While important, Text-to-CAD has received limited attention. Most studies do not utilize text to control CAD generation. Instead, they explore generating CAD designs from random noise (Wu et al., 2021; Xu et al., 2022), by randomly altering components of existing CAD designs (Xu et al., 2022, 2023; Zhang et al., 2024b), or from point cloud (Khan et al., 2024a). A few studies make preliminary attempts at Text-to-CAD (Khan et al., 2024b; Li et al., 2024b). They train Transformer-based framework with ground-truth parametric sequences as supervision, termed sequential signal.
@@ 2 Related Works
On the input side, user requirements can be expressed in diverse ways. Wu et al. (2021) uses random noise as input to generate CAD models randomly. Zhang et al. (2024b), Xu et al. (2022) and Xu et al. (2023) modify specific parts of the existing CAD models to generate new ones. Khan et al. (2024a) and Ma et al. (2024) take point cloud as input to produce corresponding CAD models. In contrast, our work focuses on textual descriptions as input. Recent studies (Khan et al., 2024b; Li et al., 2024b) explore text-based input for CAD generation. Khan et al. (2024b) proposes a data annotation pipeline for synthesizing training data and a transformer-based autoregressive network. Li et al. (2024b) designs an encoder-decoder framework with a cascading contrastive strategy and CT-Mix to align text with parametric sequences. Unlike these studies, which rely solely on sequential signals, our work combines sequential and visual signals for improved performance.
On the output side, CAD models can be represented in various formats, including Constructive Solid Geometry (CSG), Boundary Representation (B-Rep) and Sketch-and-Extrude (SE). CSG constructs 3D models by combining basic primitives such as cubes, cylinders, and spheres, through Boolean operations and subtractions (Du et al., 2018; Kania et al., 2020; Yu et al., 2021, 2023). B-Rep represents 3D models using geometric elements such as vertices, edges, and faces (Jayaraman et al., 2023; Wang et al., 2022; Xu et al., 2024). SE begins with 2D sketches composed of lines, arcs, and circles, which are then extruded to form 3D models (Willis et al., 2021; Wu et al., 2021). In this work, we adopt SE as it preserves the design history of CAD models, making them more intuitive to edit.
@@
Reinforcement Learning with Human Feedback (RLHF). RLHF has been widely applied to align model output with human preference across various domains, including LLMs (Brown et al., 2020; Radford et al., 2021; Meta, 2024), text-to-image models (liang2024richhumanfeedbacktexttoimage) and text-to-video models (Wu et al., 2024). As human annotation in RLHF is costly and not easily scalable, reinforcement learning on AI feedback (RLAIF) (Liu et al., 2023; Zhang et al., 2024a; Lee et al., 2024), which leverages machine learning models to annotate data, has been proposed as a more affordable alternative to RLHF. Since RLHF/RLAIF pipeline are complex, direct preference optimization (DPO) (Rafailov et al., 2024), which directly optimize a model to adhere to human preferences, has been proposed to avoid explicit reward modeling or reinforcement learning. In this work, we adopt DPO to address the challenge of non-differentiable rendering when learning from visual signals, as it offers a simpler yet effective solution compared to RLHF. Besides, inspired by RLAIF, we propose an automatic scoring pipeline for CAD models using LVMs. The generated scores are used to construct preference data, enabling efficient learning without reliance on costly human annotations.
@@ 3.2 Sequential Learning Stage
2) CAD Parametric Sequence Format. We adopt the format proposed by Zhang et al. (2024b) (Figure 1(b)), which represents CAD parametric sequences as text tokens rather than binary representations or numerical attributes (Xu et al., 2022, 2023; Wu et al., 2021). This text-based format simplifies processing and interpretation by LLMs.
@@ 4.1 Setups
Datasets. For the dataset used in the sequential learning stage, we use DeepCAD dataset (Wu et al., 2021) as the source for CAD parametric sequences (specifically the version processed by Xu et al. (2022)). We construct a dataset compromising 20k pairs of textual instructions and CAD parametric sequence using the techniques introduced in Section 3.2 and Appendix A.3. For the preference data used in the visual feedback stage, we employ llava-onevision-qwen2-7b (Li et al., 2024a) to construct it using the method introduced in Section 3.3. For each iteration of the visual feedback, we generate approximately 1,500 preference pairs, by using 1,000 text prompts as input, sampling 5 times per prompt, and filtering out invalid or low-quality samples. For the test set, we construct it by splitting the dataset used in sequential learning into train, validation, and test sets with a 90:5:5 ratio.
Implementation Details. LLaMA-3-8b-Instruct is used as the LLM backbone, with a maximum token length of 1024. For efficient fine-tuning, we adopt Low-Rank Adaptation (LoRA) (Hu et al., 2021) with hyperparameters and . The initial sequential learning stage lasts for 40 epochs with a learning rate of , using the AdamW optimizer. Following this, we run 5 iterations of alternating visual feedback and sequential learning stages. The visual feedback stage lasts for 5 epochs on the preference data, while the sequential learning stage lasts for 1 epoch using the same dataset as the initial sequential learning stage. Training is conducted on four NVIDIA A6000-48GB SMX GPUs using PyTorch Distributed Data Parallel (DDP).
@@
Metrics. Our evaluation focuses on assessing the alignment of generated CAD models with input instructions and the overall quality of the generated CAD models. We employ the metrics at both the sequential level and visual level. First, to evaluate the correspondence between the ground-truth and generated parametric sequences, we use F1 scores following Khan et al. (2024b). Specifically, we compute F1 score for primitives (averaged over lines, arcs, and circles for brevity) and extrusions, denoted as F1-Sketch and F1-Extrusion. Second, to assess the quality of the generated CAD models, we compare the ground-truth and generated point clouds. We adopt Chamfer Distance (CD) from Khan et al. (2024b) and additional metrics from Xu et al. (2022), including Coverage (COV), which quantifies the percentage of real data covered by generated samples using CD; Minimum Matching Distance (MMD), which evaluates the closest match between generated samples and real data; and Jensen-Shannon Divergence (JSD), which measures distribution similarity. Additionally, we compute the Invalidity Ratio (IR), which quantifies the percentage of generated parametric sequences that fail to render into valid visual objects. Furthermore, we introduce an LVM-based metric, denoted as LVM Score, to assess the visual correspondence between model predictions and input instructions. To this end, we employ GPT-4o with a dedicated evaluation prompt. Further details are provided in Appendix C.1. Finally, we conduct human assessments to rank generations from different baselines, denoted as Avg. Rank. Details on this evaluation can be found in Appendix C.2.
@@ 4.2 Main Results
Quantitative Evaluation. Table 1 summarizes the quantitative results comparing CADFusion with baseline methods (see Appendix C.4 for more details). Compared to GPT-4o, CADFusion outperforms it across all metrics. This suggests that while the general model may have acquired some CAD knowledge during pre-training, explicitly optimizing for Text-to-CAD, as in our approach, is crucial for improving performance, Compared to Text2CAD, CADFusion achieves comparable or better performance on all metrics, with particular strengths in metrics reflecting the visual quality such as LVM score and Avg. Rank. This highlights the effectiveness of incorporating visual signals in our approach, as opposed to Text2CAD, which relies solely on sequential signals. This outcome also aligns with Khan et al. (2024b)’s limitation statement that Text2CAD is limited to generating only rectangular and cylindrical shapes. When faced with complex geometries, it struggles to perform effectively.
Qualitative Evaluation. Figure 5 compares the results among the ground truth, our method, GPT-4o, and Text2CAD on the test set. GPT-4o frequently fails to produce renderable results across most test cases, which aligns with its high invalidity ratio (IR) reported in Table 1. While it occasionally generates valid shapes, its outputs are often misaligned with the input prompts. Text2CAD generate well-formed shapes without irregular edges or corners. However, it often produces oversimplified shapes and, for more complex prompts, tends to generate multiple cubes or panels instead of accurately capturing the intended structure. This aligns with its low invalidity ratio (IR) but poor visual scores. such as LVM score and Avg. Rank, in Table 1. CADFusion provides the most precise response to input instructions and achieves the highest similarity to the ground truth. It successfully captures complex shapes, including rectangles, hexagons, and nested structures, such as a hexagonal hole within a cylinder. Additionally, it exhibits a strong understanding of language cues, accurately interpreting numerical and qualitative descriptors like “long" or “T-shape". Additional qualitative results, as well as our model’s ability to generate multiple varied outputs, are presented and discussed in Appendix C.6 and C.7.
@@ 4.3 Ablation Studies
Visual Feedback. To assess the importance of visual feedback, we conduct an ablation study on CADFusion using only sequential learning, denoted as CADFusion. The first row of Table 2 presents its LVM score and invalidity ratio. Compared to our approach, denoted as CADFusion in Table 2, while CADFusion improves the invalidity ratio by 1.36%, it results in a significant decrease in the LVM score. This underscores the crucial role of the visual feedback stage: by leveraging visual preference data, our framework effectively enhances the visual quality of the generated CAD models. Additionally, CADFusion outperforms the baseline method, Text2CAD, which also relies solely on sequential signals. Note that this advantage is achieved using 20k data, while Text2CAD uses 150k data. This demonstrates the effectiveness of the techniques employed in our sequential learning stage, including leveraging LLMs as the backbone, representing CAD parametric sequences as textual tokens, and utilizing human annotations (Section 3.2).
@@ 5 Limitation
CADFusion’s results are overall promising. However, there are limitations that could be addressed in future work. First, modern LVMs show a significant performance drop when handling multiple images as input. Currently, we can only provide LVM with a single-view image to ensure both accurate image understanding and prompt following. This limitation prevents us from achieving a more effective Visual Feedback pipeline and evaluator. Second, CADFusion struggles to generate very complex shapes that require spatial and commonsense reasoning, such as the shapes of letters and words (see Appendix C.8).
@@ 6 Conclusion
Appendix A Additional Dataset Construction Detail
A.1 Converting Raw Data into Strings
@@
Arc: Defined by an arc identifier and two coordinates.222While lines and arcs generally require 2 and 3 coordinates for representation, respectively, this work leverages a simplified representation where the endpoints of lines and arcs are determined by the first point of the subsequent curve. If the loop is closed at the current curve, its endpoint is determined by the first curve in the loop.
@@
A.2 Generating Textual Instructions
@@
A.3 Dataset Construction
The dataset construction process is illustrated in Figure 6. Starting with a CAD representation from the original dataset, we generate a paired textual instruction and a stringified CAD representation. The textual instruction is created through the captioning process detailed in Section A.2, while the ground truth reference is obtained by converting the CAD formatting as described in Appendix A.1.
Appendix B Additional Training Detail
B.1 Sequential Learning
@@ B.2 Visual Feedback
Appendix C Additional Experimental Results
C.1 LVM Evaluation Setups
As mentioned in the main paper, we used a GPT-4o model as the LVM evaluator. It is selected over LLaMA-ov because we attempt to prevent the impact of AI bias that makes it prefer its own generation 333We acknowledge that this choice may introduce bias in the GPT-4o results. However, we have decided to proceed with it for two reasons: 1) our primary focus is on comparing Text2CAD with our model, and 2) the GPT-based generations during our experiment showed a significant margin in LVM scores compared to other methods, so the impact of this bias is minimal.. The prompt we used for LVM evaluation is detailed in Listing 4:
C.2 Human Evaluation Setups
We generate a quadruple of outputs for each test set instruction. Each quadruple presents four rendered generations from CADFusion, CADFusion, GPT-4o and Text2CAD, respectively. The generations are tested by their correspondence between CAD shapes and instructions. Six human judges are asked to rank the generations444Due to the lack of overlapping, we obtain approximately 50 unique samples. with the first place being the best model, and one LVM is deployed to score single-view images to the scale of 10.
C.3 GPT Baselines

### v3

@@ Abstract
Creating Computer-Aided Design (CAD) models requires considerable expertise and effort. Text-to-CAD, which converts textual descriptions into CAD parametric sequences, is crucial in streamlining this process. Recent studies have utilized ground-truth parametric sequences, known as sequential signals, as supervision to achieve this goal. However, CAD models are inherently multimodal, comprising parametric sequences and corresponding rendered visual objects. Besides, the rendering process from parametric sequences to visual objects is many-to-one. Therefore, both sequential and visual signals are critical for effective training. In this work, we introduce CADFusion, a framework that uses Large Language Models (LLMs) as the backbone and alternates between two training stages: the sequential learning (SL) stage and the visual feedback (VF) stage. In the SL stage, we train LLMs using ground-truth parametric sequences, enabling the generation of logically coherent parametric sequences. In the VF stage, we reward parametric sequences that render into visually preferred objects and penalize those that do not, allowing LLMs to learn how rendered visual objects are perceived and evaluated. These two stages alternate throughout the training, ensuring balanced learning and preserving benefits of both signals. Experiments demonstrate that CADFusion improves performance, both qualitatively and quantitatively. Code is available at https://github.com/microsoft/CADFusion.
@@ 1 Introduction
While important, Text-to-CAD has received limited attention. Most studies do not utilize text to control CAD generation. Instead, they explore generating CAD designs from random noise (Wu et al., 2021; Xu et al., 2022), by randomly altering components of existing CAD designs (Xu et al., 2022, 2023; Zhang et al., 2025), or from point cloud (Khan et al., 2024a). A few studies make preliminary attempts at Text-to-CAD (Khan et al., 2024b; Li et al., 2024b). They train Transformer-based framework with ground-truth parametric sequences as supervision, termed sequential signal.
@@ 2 Related Works
On the input side, user requirements can be expressed in diverse ways. Wu et al. (2021) uses random noise as input to generate CAD models randomly. Zhang et al. (2025), Xu et al. (2022) and Xu et al. (2023) modify specific parts of the existing CAD models to generate new ones. Khan et al. (2024a) and Ma et al. (2024) take point cloud as input to produce corresponding CAD models, while Seff et al. (2022) uses hand sketches. In contrast, our work focuses on textual descriptions as input, leverages LLM as backbone on stringified representations, and does not adapt codebook encoders such as VQ-VAE to generate initial representations. Recent studies (Khan et al., 2024b; Li et al., 2024b) explore text-based input for CAD generation. Khan et al. (2024b) proposes a data annotation pipeline for synthesizing training data and a transformer-based autoregressive network. Li et al. (2024b) designs an encoder-decoder framework with a cascading contrastive strategy and CT-Mix to align text with parametric sequences. Unlike these studies, which rely solely on sequential signals, our work combines sequential and visual signals for improved performance.
On the output side, CAD models can be represented in various formats, including Constructive Solid Geometry (CSG), Boundary Representation (B-Rep) and Sketch-and-Extrude (SE). CSG constructs 3D models by combining basic primitives such as cubes, cylinders, and spheres, through Boolean operations and subtractions (Du et al., 2018; Kania et al., 2020; Yu et al., 2022, 2024). B-Rep represents 3D models using geometric elements such as vertices, edges, and faces (Jayaraman et al., 2023; Wang et al., 2022; Xu et al., 2024). SE begins with 2D sketches composed of lines, arcs, and circles, which are then extruded to form 3D models (Willis et al., 2021; Wu et al., 2021). In this work, we adopt SE as it preserves the design history of CAD models, making them more intuitive to edit.
@@
Reinforcement Learning with Human Feedback (RLHF). RLHF has been widely applied to align model output with human preference across various domains, including LLMs (Brown et al., 2020; Radford et al., 2021; Meta, 2024), text-to-image models (Liang et al., 2024) and text-to-video models (Wu et al., 2024). As human annotation in RLHF is costly and not easily scalable, reinforcement learning on AI feedback (RLAIF) (Liu et al., 2023; Zhang et al., 2024; Lee et al., 2024), which leverages machine learning models to annotate data, has been proposed as a more affordable alternative to RLHF. Since RLHF/RLAIF pipeline are complex, direct preference optimization (DPO) (Rafailov et al., 2024), which directly optimize a model to adhere to human preferences, has been proposed to avoid explicit reward modeling or reinforcement learning. In this work, we adopt DPO to address the challenge of non-differentiable rendering when learning from visual signals, as it offers a simpler yet effective solution compared to RLHF. Besides, inspired by RLAIF, we propose an automatic scoring pipeline for CAD models using LVMs. The generated scores are used to construct preference data, enabling efficient learning without reliance on costly human annotations.
@@ 3.2 Sequential Learning Stage
2) CAD Parametric Sequence Format. We adopt the format proposed by Zhang et al. (2025) (Figure 1(b)), which represents CAD parametric sequences as text tokens rather than binary representations or numerical attributes (Xu et al., 2022, 2023; Wu et al., 2021). This text-based format simplifies processing and interpretation by LLMs.
@@ 4.1 Setups
Datasets. For the dataset used in the sequential learning stage, we use DeepCAD dataset (Wu et al., 2021) as the source for CAD parametric sequences (specifically the version processed by Xu et al. (2022)). We construct a dataset compromising 20k pairs of textual instructions and CAD parametric sequence using the techniques introduced in Section 3.2 and Appendix B.3. For the preference data used in the visual feedback stage, we employ llava-onevision-qwen2-7b (Li et al., 2024a) to construct it using the method introduced in Section 3.3. For each iteration of the visual feedback, we generate approximately 1,500 preference pairs, by using 1,000 text prompts as input, sampling 5 times per prompt, and filtering out invalid or low-quality samples. For the test set, we construct it by splitting the dataset used in sequential learning into train, validation, and test sets with a 90:5:5 ratio.
Implementation Details. LLaMA-3-8b-Instruct is used as the LLM backbone, with a maximum token length of 1024. For efficient fine-tuning, we adopt Low-Rank Adaptation (LoRA) (Hu et al., 2022) with hyperparameters and . The initial sequential learning stage lasts for 40 epochs with a learning rate of , using the AdamW optimizer. Following this, we run 5 iterations of alternating visual feedback and sequential learning stages. The visual feedback stage lasts for 5 epochs on the preference data, while the sequential learning stage lasts for 1 epoch using the same dataset as the initial sequential learning stage. Training is conducted on four NVIDIA A6000-48GB SMX GPUs using PyTorch Distributed Data Parallel (DDP).
@@
Metrics. Our evaluation focuses on assessing the alignment of generated CAD models with input instructions and the overall quality of the generated CAD models. We employ the metrics at both the sequential level and visual level. First, to evaluate the correspondence between the ground-truth and generated parametric sequences, we use F1 scores following Khan et al. (2024b). Specifically, we compute F1 score for primitives (averaged over lines, arcs, and circles for brevity) and extrusions, denoted as F1-Sketch and F1-Extrusion. Second, to assess the quality of the generated CAD models, we compare the ground-truth and generated point clouds. We adopt Chamfer Distance (CD) from Khan et al. (2024b) and additional metrics from Xu et al. (2022), including Coverage (COV), which quantifies the percentage of real data covered by generated samples using CD; Minimum Matching Distance (MMD), which evaluates the closest match between generated samples and real data; and Jensen-Shannon Divergence (JSD), which measures distribution similarity. Additionally, we compute the Invalidity Ratio (IR), which quantifies the percentage of generated parametric sequences that fail to render into valid visual objects. Furthermore, we introduce an LVM-based metric, denoted as LVM Score, to assess the visual correspondence between model predictions and input instructions. To this end, we employ GPT-4o with a dedicated evaluation prompt. Further details are provided in Appendix D.1. Finally, we conduct human assessments to rank generations from different baselines, denoted as Avg. Rank. Details on this evaluation can be found in Appendix D.2.
@@ 4.2 Main Results
Quantitative Evaluation. Table 1 summarizes the quantitative results comparing CADFusion with baseline methods (see Appendix D.4 for more details). Compared to GPT-4o, CADFusion outperforms it across all metrics. This suggests that while the general model may have acquired some CAD knowledge during pre-training, explicitly optimizing for Text-to-CAD, as in our approach, is crucial for improving performance, Compared to Text2CAD222Our comparison with Text2CAD is not entirely aligned and is in favor of it. Performing poorly on prompts we provided, we have present the results of Text2CAD tested with their original prompts. We detail this problem in Appendix D.4., CADFusion achieves comparable or better performance on all metrics, with particular strengths in metrics reflecting the visual quality such as LVM score and Avg. Rank. This highlights the effectiveness of incorporating visual signals in our approach, as opposed to Text2CAD, which relies solely on sequential signals. This outcome also aligns with Khan et al. (2024b)’s limitation statement that Text2CAD is limited to generating only rectangular and cylindrical shapes. When faced with complex geometries, it struggles to perform effectively.
Qualitative Evaluation. Figure 5 compares the results among the ground truth, our method, GPT-4o, and Text2CAD on the test set. GPT-4o frequently fails to produce renderable results across most test cases, which aligns with its high invalidity ratio (IR) reported in Table 1. While it occasionally generates valid shapes, its outputs are often misaligned with the input prompts. Text2CAD generate well-formed shapes without irregular edges or corners. However, it often produces oversimplified shapes and, for more complex prompts, tends to generate multiple cubes or panels instead of accurately capturing the intended structure. This aligns with its low invalidity ratio (IR) but poor visual scores. such as LVM score and Avg. Rank, in Table 1. CADFusion provides the most precise response to input instructions and achieves the highest similarity to the ground truth. It successfully captures complex shapes, including rectangles, hexagons, and nested structures, such as a hexagonal hole within a cylinder. Additionally, it exhibits a strong understanding of language cues, accurately interpreting numerical and qualitative descriptors like “long" or “T-shape". Additional qualitative results, as well as our model’s ability to generate multiple varied outputs, are presented and discussed in Appendix D.6 and D.7.
@@ 4.3 Ablation Studies
Visual Feedback. To assess the importance of visual feedback, we conduct an ablation study on CADFusion using only sequential learning, denoted as CADFusion. The first row of Table 2 presents its LVM score and invalidity ratio. Compared to our approach, denoted as CADFusion in Table 2, while CADFusion improves the invalidity ratio by 1.36%, it results in a considerable decrease in the LVM score. This underscores the crucial role of the visual feedback stage: by leveraging visual preference data, our framework effectively enhances the visual quality of the generated CAD models. Additionally, CADFusion outperforms the baseline method, Text2CAD, which also relies solely on sequential signals. Note that this advantage is achieved using 20k data, while Text2CAD uses 150k data. This demonstrates the effectiveness of the techniques employed in our sequential learning stage, including leveraging LLMs as the backbone, representing CAD parametric sequences as textual tokens, and utilizing human annotations (Section 3.2).
@@ 5 Limitation
CADFusion’s results are overall promising. However, there are limitations that could be addressed in future work. First, modern LVMs suffer from performance drop when handling multiple images as input. Currently, we can only provide LVM with a single-view image to ensure both accurate image understanding and prompt following. This limitation prevents us from achieving a more effective Visual Feedback pipeline and evaluator. Second, CADFusion struggles to generate very complex shapes that require spatial and commonsense reasoning, such as the shapes of letters and words (see Appendix D.8).
@@ 6 Conclusion
Acknowledgement
We thank Weijian Ma for insightful conversation and discussion. We also appreciate the informative comments and suggestions provided by the anonymous reviewers.
@@ Acknowledgement
Appendix A User Guidelines for Prompting
A good prompt follows a structured description: (1) shape overview, (2) shape details, and (3) shape applications. Given the varying shape complexity, we encourage but do not enforce describing item (2) and (3). Below is an example caption retrieved from Figure 8, Row 2, Item 3, demonstrating this approach:
Appendix B Additional Dataset Construction Detail
B.1 Converting Raw Data into Strings
@@
Arc: Defined by an arc identifier and two coordinates.333While lines and arcs generally require 2 and 3 coordinates for representation, respectively, this work leverages a simplified representation where the endpoints of lines and arcs are determined by the first point of the subsequent curve. If the loop is closed at the current curve, its endpoint is determined by the first curve in the loop.
@@
B.2 Generating Textual Instructions
@@
B.3 Dataset Construction
The dataset construction process is illustrated in Figure 6. Starting with a CAD representation from the original dataset, we generate a paired textual instruction and a stringified CAD representation. The textual instruction is created through the captioning process detailed in Section B.2, while the ground truth reference is obtained by converting the CAD formatting as described in Appendix B.1.
Appendix C Additional Training Detail
C.1 Sequential Learning
@@ C.2 Visual Feedback
Appendix D Additional Experimental Results
D.1 LVM Evaluation Setups
As mentioned in the main paper, we used a GPT-4o model as the LVM evaluator. It is selected over LLaMA-ov because we attempt to prevent the impact of AI bias that makes it prefer its own generation 444We acknowledge that this choice may introduce bias in the GPT-4o results. However, we have decided to proceed with it for two reasons: 1) our primary focus is on comparing Text2CAD with our model, and 2) the GPT-based generations during our experiment showed a big margin in LVM scores compared to other methods, so the impact of this bias is minimal.. The prompt we used for LVM evaluation is detailed in Listing 4:
D.2 Human Evaluation Setups
We generate a quadruple of outputs for each test set instruction. Each quadruple presents four rendered generations from CADFusion, CADFusion, GPT-4o and Text2CAD, respectively. The generations are tested by their correspondence between CAD shapes and instructions. Six human judges (with college-level or higher education records) are asked to rank the generations555Due to the lack of overlapping, we obtain approximately 50 unique samples. with the first place being the best model. The ranks are then collected and averaged to be the scores we presented as human evaluation.
D.3 GPT Baselines